# Avoiding Catastrophe in Online Learning by Asking for Help

### v3

@@ 5.2 Proof sketch
Let $U$ be the uniform distribution over $S$. For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{s\sim U}[\pi(s)\neq\tilde{\pi}(s)]\leq\varepsilon$.
@@
A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\mathbf{s}\in S^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(s_{t})=\tilde{\pi}(s_{t})$ for all $t\in[T]$.
@@
Lemma 5.1 (Lemma 7.3.2 in [16] [^11] See also [19] or Lemma 13.6 in [6] for variants which are less convenient for our purposes.).
@@
Lemma 5.2 (Lemmas 21.13 and A.5 in [34]).
@@
Lemma 5.3 (Equation 2 and Lemma 3.3 in [17]).
Let $\tilde{\Pi}$ be a finite smooth $\varepsilon$-cover of $\Pi$ and let $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ be a sequence of $\sigma$-smooth distributions. If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then
$\operatorname*{\mathbb{E}}_{\mathbf{s}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(s_{t})\neq\tilde{\pi}(s_{t}))\right]\in O\left(\sqrt{\frac{d\varepsilon T}{\sigma}\log T\log(1/\varepsilon)}+\frac{\varepsilon T}{\sigma}\log T\right)$
We will run a variant of Hedge on $\tilde{\Pi}$. The vanilla Hedge algorithm operates in the standard online learning model where on each time step, the agent selects a policy (or more generally, a hypothesis), and observes the loss of every policy. In general the loss function can depend arbitrarily on the time step, the policy, and prior events, but we will only use the indicator loss function $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$. Crucially, whenever we query and learn $\pi^{m}(s_{t})$, we can compute $\ell(t,\pi)$ for every $\pi\in\tilde{\Pi}$.
We cannot afford to query on every time step, however. Recently, [33] analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”. They proved a regret bound when a query is issued on each time step with fixed probability $p$. Lemma 5.4 restates their result in a form that is more convenient for us (see Appendix B for details). Although their result is stated for non-adaptive adversaries, we explain in Section B.3 how their argument easily generalizes to adaptive adversaries. Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
Lemma 5.4 (Lemma 3.5 in [33]).
@@
We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(s_{t})\neq\pi^{m}(s_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case). This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
The other key ingredient of the proof is analyzing the “ask for help when out-of-distribution” component. Combined with the local generalization assumption, this allows us to fairly easily bound the cost of a single mistake (Lemma B.2). The trickier part is bounding the number of resulting queries. It is tempting to claim that the states queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true. Instead, the packing analysis must be performed on a per-action basis, and we reuse Lemma B.1 to crucially show that there isn’t too much “spillover” between actions. The resulting overall query bound is given by Lemma B.7.
@@ 6 Conclusion and future work
We are also interested in alternatives to the local generalization assumption, since an action which is safe in one state may not always be safe in a nearby state (depending on the definition of “nearby”). Some assumption is necessary, or there is no way to avoid costly mistakes without querying on every time step. One possibility is Bayesian inference. We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
@@
Acknowledgements
This work was supported in part by a gift from Open Philanthropy to the Center for Human-Compatible AI (CHAI) at UC Berkeley. This paper also benefited from discussions with many other researchers. We would like to especially thank (in alphabetical order) Cameron Allen, Michael Cohen, Scott Emmons, Nika Haghtalab, Aly Lidayan, Juan Liévano-Karim, Bhaskar Mishra, and Matteo Russo.
@@ A.1 Proof roadmap
Recall that $\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\mathbf{a}}$ denotes the expectation over both the random states and any additional randomness in the algorithm.
The proof proceeds via the following steps:
@@
    Provide a simple variant of the Chernoff bound which we will apply multiple times (Lemma A.2).
@@
    Show that with high probability, $\sum_{j\in A}|V_{j}|$ is adequately large (Lemma A.3).
@@
    The key lemma is Lemma A.4, which shows that a randomly sampled $\mathbf{x}$ produces poor agent performance with high probability. The central idea is that at least $f(T)-|Q_{T}|$ sections are never queried (which is large, by Lemma A.1), so the agent has no way of knowing the optimal action in those sections. As a result, the agent picks the wrong answer at least half the time on average (and at least a quarter of the time with high probability). Lemma A.3 implies that a constant fraction of those time steps will have quite suboptimal payoffs, again with high probability.
@@
    To complete the proof, we observe that $\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$. This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\mathbf{x}}$ has high expected regret, then the worst case $\mu$ also has high expected regret.
Note that $\mathbf{s},\mathbf{a}$, and $\mathbf{x}$ are random variables, so all variables defined on top of them (such as $V_{j}$) are also random variables. In contrast, the partition $S=\{S_{1},\dots,S_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
Lastly, while the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different. However, bounds of the form discussed in Section 4.1 can still be found as an intermediate step in Part 4 of the proof of Lemma A.4.
@@ A.2 Proof
Let $z_{1},\dots,z_{n}$ be i.i.d. variables in $\{0,1\}$ and let $Z=\sum_{i=1}^{n}z_{i}$. If $\operatorname*{\mathbb{E}}[Z]\geq M$, then $\Pr\big{[}Z\leq M/2\big{]}\leq\exp(-M/8)$.
@@
By the Chernoff bound for i.i.d. binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$. Since $-\operatorname*{\mathbb{E}}[Z]\leq-M$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-M/8)$. Also, $M/2\leq E[Z]/2$ implies $\Pr[Z\leq M/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$. Combining these inequalities proves the lemma. ∎
@@
$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
@@
$\Pr\left[\sum_{j\in A}|V_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
@@
Independently sample $\mathbf{x}\sim U(\{0,1\}^{f(T)})$ and $\mathbf{s}\sim U^{T}$. [^12] That is, the entire set $\{x_{1},\dots,x_{f(T)},s_{1},\dots,s_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$,
$\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
@@
Part 1: setup. Let $J_{\neg Q}=\{j\in[f(T)]:s_{t}\not\in S_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried. Since each query appears in exactly one section (because each state appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|$.
For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $V_{j}$:
$y_{j}=\operatorname*{arg\,max}_{a\in\{0,1\}}\Big{|}\{t\in V_{j}:a=a_{t}\}\Big{|}$
Let $\bar{J}=\{j\in J_{\neg Q}:x_{j}\neq y_{j}\}$. For each $j\in\bar{J}$, let $V_{j}^{\prime}=\{t\in V_{j}:a_{t}\neq x_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\mathbf{x}}$).
@@
$\Pr\left[|\bar{J}|\leq\frac{f(T)-|Q_{T}|}{4}\right]\leq\exp\left(-\frac{f(T)-|Q_{T}|}{16}\right)$
Part 3: $|V_{j}^{\prime}|\geq|V_{j}|/2$. Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in V_{j}$, so $a_{t}\in\{0,1\}$ for all $t\in V_{j}$. Since the agent chooses one of two actions for each $t\in V_{j}$, the more frequent action must be chosen chosen at least half of the time: $a_{t}=y_{j}$ for at least half of the time steps in $V_{j}$. Since $x_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $a_{t}=y_{j}\neq x_{j}$ for those time steps, so $|V_{j}^{\prime}|\geq|V_{j}|/2$.
Part 4: a bound in terms of $\bar{J}$ and $V_{j}$. Consider any $j\in\bar{J}$ and $t\in V_{j}^{\prime}\subseteq V_{j}$. By definition of $V_{j}$, we have $|m_{j}-s_{t}|\leq\frac{1}{4f(T)}$. Then by definition of $\mu_{f,\mathbf{x}}$,
$\mu_{f,\mathbf{x}}(s_{t},a_{t})=$ $\ 1-L\left(\frac{1}{2f(T)}-|s_{t}-m_{j}|\right)$
@@
$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|V_{j}|/2}$
$=$ $\ \sum_{j\in\bar{J}}\frac{|V_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
@@
$\log\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|V_{j}|}{8f(T)}$
@@
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-|Q_{T}|}{4}$
$=$ $\ -\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}$
@@
Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-|Q_{T}|}{16}\big{)}$ for brevity. See 4.1
@@
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ for all $\mathbf{s},\mu,\pi^{m}$. Let $f(T)=\sqrt{(g(T)+1)T}$. Then by Lemma A.1, $g(T)\in o(f(T))$ and $f(T)\in o(T)$. Combining this with $|Q_{T}|\leq g(T)$, we get $\lim_{T\to\infty}\alpha(T)=0$. Also, since $|Q_{T}|\in o(f(T))$, there exists $T_{0}$ such that $|Q_{T}|\leq f(T)/2$ for all $T\geq T_{0}$. Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\leq 1$, we have
$\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ \>\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})$
$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|Q_{T}|)}{2^{7}f(T)^{2}}\right)$
@@
whenever $T\geq T_{0}$. Since $\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})=1$ always, we have [^13] Fubini’s theorem means we need not worry about the order of the expectation operators.
$\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu_{f,\mathbf{x}},\pi^{m})$
$=\operatorname*{\mathbb{E}}_{\mathbf{x}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\mathbf{x}}^{m}(s_{t})-\prod_{t=1}^{T}\mu_{f,\mathbf{x}}(s_{t},a_{t})\right]$
@@
$\lim_{T\to\infty}\sup_{\mu}\ \operatorname*{\mathbb{E}}_{\mathbf{s}\sim U^{T},\,\mathbf{a}}\ R_{T}(\mathbf{s},\mathbf{a},\mu,\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$

### v4

@@ 5.2 Proof sketch
Let $U$ be the uniform distribution over $\mathcal{X}$. For $\varepsilon>0$, a policy class $\tilde{\Pi}$ is a smooth $\varepsilon$-cover of a policy class $\Pi$ is for every $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\Pr_{x\sim U}[\pi(x)\neq\tilde{\pi}(x)]\leq\varepsilon$.
@@
A policy class $\tilde{\Pi}$ is an adversarial cover of a policy class $\Pi$ is for every $\boldsymbol{x}\in\mathcal{X}^{T}$ and $\pi\in\Pi$, there exists $\tilde{\pi}\in\tilde{\Pi}$ such that $\pi(x_{t})=\tilde{\pi}(x_{t})$ for all $t\in[T]$.
@@
Lemma 5.1 (Lemma 7.3.2 in Haghtalab (2018) [^13] See also Haussler & Long (1995) or Lemma 13.6 in Boucheron et al. (2013) for variants which are less convenient for our purposes.).
@@
Lemma 5.2 (Lemmas 21.13 and A.5 in Shalev-Shwartz & Ben-David (2014)).
@@
Lemma 5.3 (Equation 2 and Lemma 3.3 in Haghtalab et al. (2024)).
Let $\tilde{\Pi}$ be a finite smooth $\varepsilon$-cover of $\Pi$ and let $\boldsymbol{\mathcal{D}}=\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ be a sequence of $\sigma$-smooth distributions. If $\varepsilon T\log T>12\sigma d\log(4e^{2}/\varepsilon)$, then $\operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim\boldsymbol{\mathcal{D}}}\left[\sup\limits_{\pi\in\Pi}\min\limits_{\tilde{\pi}\in\tilde{\Pi}}\sum_{t=1}^{T}\mathbf{1}(\pi(x_{t})\neq\tilde{\pi}(x_{t}))\right]\in O\left(\frac{1}{\sigma}T\varepsilon\log T\sqrt{d\log(1/\varepsilon)}\right)$.
We will run a variant of Hedge on $\tilde{\Pi}$. The vanilla Hedge algorithm operates in the standard online learning model where on each time step, the agent selects a policy (or more generally, a hypothesis), and observes the loss of every policy. In general the loss function can depend arbitrarily on the time step, the policy, and prior events, but we will only use the indicator loss function $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$. Crucially, whenever we query and learn $\pi^{m}(x_{t})$, we can compute $\ell(t,\pi)$ for every $\pi\in\tilde{\Pi}$.
We cannot afford to query on every time step, however. Recently, Russo et al. (2024) analyzed a variant of Hedge where losses are observed only in response to queries, which they call “label-efficient feedback”. They proved a regret bound when a query is issued on each time step with fixed probability $p$. Lemma 5.4 restates their result in a form that is more convenient for us (see Appendix B for details). Although their result is stated for non-adaptive adversaries, we explain in Section B.3 how their argument easily generalizes to adaptive adversaries. Full pseudocode for HedgeWithQueries can also be found in the appendix (Algorithm 2).
Lemma 5.4 (Lemma 3.5 in Russo et al. (2024)).
@@
We apply Lemma 5.4 with $\ell(t,\pi)=\mathbf{1}(\pi(x_{t})\neq\pi^{m}(x_{t}))$ and combine this with Lemmas 5.1 and 5.3 (in the $\sigma$-smooth case) and with Lemma 5.2 (in the adversarial case). This yields a $O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$ bound on the number of mistakes made by Algorithm 1 (Lemma B.1).
The other key ingredient of the proof is analyzing the “ask for help when out-of-distribution” component. Combined with the local generalization assumption, this allows us to fairly easily bound the cost of a single mistake (Lemma B.2). The trickier part is bounding the number of resulting queries. It is tempting to claim that the inputs queried in the out-of-distribution case must all be separated by at least $\varepsilon^{1/n}$ and thus form an $\varepsilon^{1/n}$-packing, but this is actually not true. Instead, we provide a novel method for bounding the number of data points (i.e., queries) needed to cover a set with respect to the realized actions of the algorithm (Lemma B.7). This is in contrast to vanilla packing arguments which consider all data points in aggregate. Our method may be useful in other contexts where a more refined packing argument is needed.
@@ 6 Conclusion and future work
We are also interested in alternatives to the local generalization assumption. We should expect some assumption to be necessary: if not, the payoff function $\mu(x,y)=\mathbf{1}(\pi^{m}(x)=y)$ means that the agent must make zero mistakes (with high probability), which turns out to be impossible even for $\sigma$-smooth $\boldsymbol{x}$ and finite VC dimension (Theorem E.3). One possible alternative is Bayesian inference. We intentionally avoided Bayesian approaches in this paper due to tractability concerns, but it seems premature to abandon those ideas entirely.
@@ A.1 Proof roadmap
The proof considers an arbitrary algorithm with sublinear queries, and proceeds via the following steps:
@@
    Prove a simple variant of the Chernoff bound which we will apply multiple times (Lemma A.2).
@@
    Show that with high probability, $\sum_{j\in A}|M_{j}|$ is adequately large (Lemma A.3).
@@
    The key lemma is Lemma A.4, which shows that a randomly sampled $\boldsymbol{a}$ produces poor agent performance with high probability. The central idea is that at least $f(T)-|Q_{T}|$ sections are never queried (which is large, by Lemma A.1), so the agent has no way of knowing the optimal action in those sections. As a result, the agent picks the wrong answer at least half the time on average (and at least a quarter of the time with high probability). Lemma A.3 implies that a constant fraction of those time steps will have quite suboptimal payoffs, again with high probability.
@@
    Finally, $\sup\limits_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq\operatorname*{\mathbb{E}}\limits_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \operatorname*{\mathbb{E}}\limits_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\mu_{f,\boldsymbol{a}},\pi^{m})$, where $U(\{0,1\}^{f(T)})$ is the uniform distribution over bit strings of length $f(T)$. This is essentially an application of the probabilistic method: if a randomly chosen $\mu_{f,\boldsymbol{a}}$ has high expected regret, then the worst case $\boldsymbol{\mu}$ also has high expected regret.
While reading the proof, one should keep in mind the following:
  1. 1.
    Since $\boldsymbol{x},\boldsymbol{y}$, and $\boldsymbol{a}$ are random variables, all variables defined on top of them (such as $M_{j}$) are also random variables. In contrast, the partition $\mathcal{X}=\{X_{1},\dots,X_{f(T)}\}$ and properties thereof (like the midpoints $m_{j}$) are not random variables.
  2. 2.
    We will assume that the worst-case (not expected) number of queries is sublinear in $T$. Proposition E.2 shows that relaxing this to expected sublinear queries would not affect our conclusion.
  3. 3.
    While the intuition provided in Section 4.1 is accurate, the analysis will mostly occur in log space, so the bounds will look different. However, bounds of the form discussed in Section 4.1 can still be found as an intermediate step in Part 4 of the proof of Lemma A.4.
@@ A.2 Proof
Let $z_{1},\dots,z_{n}$ be i.i.d. variables in $\{0,1\}$ and let $Z=\sum_{i=1}^{n}z_{i}$. If $\operatorname*{\mathbb{E}}[Z]\geq W$, then $\Pr\big{[}Z\leq W/2\big{]}\leq\exp(-W/8)$.
@@
By the Chernoff bound for i.i.d. binary variables, we have $\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]\leq\exp(-\operatorname*{\mathbb{E}}[Z]/8)$. Since $-\operatorname*{\mathbb{E}}[Z]\leq-W$ and $\exp$ is an increasing function, we have $\exp(-\operatorname*{\mathbb{E}}[Z]/8)\leq\exp(-W/8)$. Also, $W/2\leq E[Z]/2$ implies $\Pr[Z\leq W/2]\leq\Pr[Z\leq\operatorname*{\mathbb{E}}[Z]/2]$. Combining these inequalities proves the lemma. ∎
@@
$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T}{16f(T)}\right)$
@@
$\Pr\left[\sum_{j\in A}|M_{j}|\leq\frac{T|A|}{4f(T)}\right]\leq\exp\left(\frac{-T|A|}{16f(T)}\right)\leq\exp\left(\frac{-T}{16f(T)}\right)$
@@
Let $f,g:\mathbb{N}\to\mathbb{N}$. Suppose $|Q_{T}|\leq g(T)$ always and independently sample $\boldsymbol{a}\sim U(\{0,1\}^{f(T)})$ and $\boldsymbol{x}\sim U^{T}$. [^14] That is, the entire set $\{a_{1},\dots,a_{f(T)},x_{1},\dots,x_{T}\}$ is mutually independent. Then with probability at least $1-\exp\big{(}\frac{-T}{16f(T)}\big{)}-\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$,
$\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq\exp\left(-\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}\right)$
@@
Part 1: setup. Let $J_{\neg Q}=\{j\in[f(T)]:x_{t}\not\in X_{j}\ \forall t\in Q_{T}\}$ be the set of sections that are never queried. Since each query appears in exactly one section (because each input appears in exactly one section), $|J_{\neg Q}|\geq f(T)-|Q_{T}|\geq f(T)-g(T)$.
For each $j\in J_{\neg Q}$, let $y_{j}$ be the action taken most frequently among time steps in $M_{j}$:
$y_{j}=\operatorname*{arg\,max}_{y\in\{0,1\}}\Big{|}\{t\in M_{j}:y=y_{t}\}\Big{|}$
Let $\bar{J}=\{j\in J_{\neg Q}:a_{j}\neq y_{j}\}$. For each $j\in\bar{J}$, let $M_{j}^{\prime}=\{t\in M_{j}:y_{t}\neq a_{j}\}$ be the set of time steps where the agent chooses the wrong action (assuming payoff function $\mu_{f,\boldsymbol{a}}$).
@@
$\Pr\left[|\bar{J}|\leq\frac{f(T)-g(T)}{4}\right]\leq\exp\left(-\frac{f(T)-g(T)}{16}\right)$
Part 3: $|M_{j}^{\prime}|\geq|M_{j}|/2$. Since $j\in J_{\neg Q}$, the mentor is not queried on any time step $t\in M_{j}$, so $y_{t}\in\{0,1\}$ for all $t\in M_{j}$. Since the agent chooses one of two actions for each $t\in M_{j}$, the more frequent action must be chosen chosen at least half of the time: $y_{t}=y_{j}$ for at least half of the time steps in $M_{j}$. Since $a_{j}\neq y_{j}$ for $j\in\bar{J}$, we have $y_{t}=y_{j}\neq a_{j}$ for those time steps, so $|M_{j}^{\prime}|\geq|M_{j}|/2$.
Part 4: a bound in terms of $\bar{J}$ and $M_{j}$. Consider any $j\in\bar{J}$ and $t\in M_{j}^{\prime}\subseteq M_{j}$. By definition of $M_{j}$, we have $|m_{j}-x_{t}|\leq\frac{1}{4f(T)}$. Then by definition of $\mu_{f,\boldsymbol{a}}$,
$\mu_{f,\boldsymbol{a}}(x_{t},y_{t})=$ $\ 1-L\left(\frac{1}{2f(T)}-|x_{t}-m_{j}|\right)$
@@
$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ \log\prod_{j\in\bar{J}}\left(1-\frac{L}{4f(T)}\right)^{|M_{j}|/2}$
$=$ $\ \sum_{j\in\bar{J}}\frac{|M_{j}|}{2}\log\left(1-\frac{L}{4f(T)}\right)$
$\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
@@
$\log\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq$ $\ -\sum_{j\in\bar{J}}\frac{L|M_{j}|}{8f(T)}$
@@
$\leq$ $\ -\frac{L}{8f(T)}\cdot\frac{T}{4f(T)}\cdot\frac{f(T)-g(T)}{4}$
$=$ $\ -\frac{LT(f(T)-g(T))}{2^{7}f(T)^{2}}$
@@
Let $\alpha(T)=\exp\big{(}\frac{-T}{16f(T)}\big{)}+\exp\big{(}-\frac{f(T)-g(T)}{16}\big{)}$ for brevity (with $f$ and $g$ to be chosen later). See 4.1
@@
If the algorithm has sublinear queries, then there exists $g(T)\in o(T)$ such that $|Q_{T}|\leq g(T)$ always. Then by Lemma A.1, $g(T)\in o(\sqrt{g(T)T})$ and $\sqrt{g(T)T}\in o(T)$. Let $f(T)=\lceil\sqrt{g(T)T}\rceil$: then $f(T)\in\Theta(\sqrt{g(T)T})$, so $g(T)\in o(f(T))$ and $f(T)\in o(T)$. First, this implies that $\lim_{T\to\infty}\alpha(T)=0$. Next, $g(T)\in o(f(T))$ implies that exists $T_{0}$ such that $g(T)\leq f(T)/2$ for all $T\geq T_{0}$. Combining this with Lemma A.4 and noting that $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\leq 1$, for all $T\geq T_{0}$ we have
$\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ \>\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})$
$\leq$ $\ \alpha(T)\cdot 1+\big{(}1-\alpha(T)\big{)}\exp\left(-\frac{LT(f(T)-|g(T))}{2^{7}f(T)^{2}}\right)$
@@
Since $\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})=1$ always, we have [^15] Fubini’s theorem means we need not worry about the order of the expectation operators.
$\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ \operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},(\mu_{f,\boldsymbol{a}},\dots,\mu_{f,\boldsymbol{a}}),\pi^{m})$
$=\operatorname*{\mathbb{E}}_{\boldsymbol{a}\sim U(\{0,1\}^{f(T)})}\ \,\operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}$ $\ \,\left[\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{f,\boldsymbol{a}}(x_{t},y_{t})\right]$
@@
$\lim_{T\to\infty}\sup_{\boldsymbol{\mu}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{x}\sim U^{T},\,\boldsymbol{y}}\ R_{T}(\boldsymbol{x},\boldsymbol{y},\boldsymbol{\mu},\pi^{m})\geq$ $\ 1-\lim_{T\to\infty}\alpha(T)-\big{(}1-\lim_{T\to\infty}\alpha(T)\big{)}\cdot\exp\left(\lim_{T\to\infty}-\frac{LT}{2^{8}f(T)}\right)$